OpenStudy (anonymous):

log (base 6) (2x + 3) = 3

jimthompson5910 (jim_thompson5910):

\[\Large \log_{6}(2x+3) = 3\] \[\Large 2x+3 = 6^3\] \[\Large 2x+3 = 216\] \[\Large 2x = 216-3\] \[\Large 2x = 213\] \[\Large x = \frac{213}{2}\]